Air that initially occupies 0.144 m3 at a gauge pressure of 211.0 kPa is expanded isothermally to a pressure of 101.3 kPa and then cooled at constant pressure until it reaches its initial volume. Compute the work done by the air. (Gauge pressure is the difference between the actual pressure and atmospheric pressure.)
Answer #1

Using Gas equation

PV = nRT

For isothermal process

P_{1}V_{1} = P_{2}V_{2}

V_{2} = \frac{P_{1}V_{1} }{P_{2}}

V_{2} = \frac{211.0 kPa\times 0.144 m^{3} }{101.3 kPa} = 0.29994 m^{3}

Work done during isothermal expansion

W_{1} = P_{1}V_{1}\ln \left ( \frac{P_{1}}{P_{2}} \right )

W_{1} = 211.0\times 10^{3}Pa\times 0.144 m^{3}\times \ln \left ( \frac{211.0}{101.3} \right )= 22294.92 J

Now work done during isobaric process

W_{2} = P_{2}(V_{1}-V_{2}) = 101.3\times 10^{3} Pa\times (0.144 m^{3}-0.29994 m^{3})= -15796.72 J

Total Work done

W = W_{1}+W_{2} =22294.92 J -15796.72 J = 6498.2 J